OpenStudy (anonymous):

1. 1-cos^2 theta/sin theta @shamil98

OpenStudy (shamil98):

Yeah this is what I was saying , I don't remember how to do this stuff. Sorry.

OpenStudy (anonymous):

No worries, thank you anyways!

OpenStudy (anonymous):

1- cos^2 = sin^2 ------ ----- = sin sin sin

OpenStudy (anonymous):

don't get what I mean? numerator is 1- cos^2 which is = \(sin^2 \theta\) , then divided by sin \(\theta\) = sin\(\theta\)

OpenStudy (anonymous):

That makes sense! Thanks. What about sin^2x+tan^2x+cos^2x/sec x ??

OpenStudy (anonymous):

sin^2 + cos^2 =1 so numerator = tan^2 + 1 = sec^2 then divided by sec = sec

OpenStudy (anonymous):

sec^2x-tan^2x/cos+sin^2x

OpenStudy (anonymous):

sec^2 = tan ^2 +1 so, numerator = tan^2 +1 - tan^2 =1 we can do nothing with the denominator( if it is as you post) . so the answer is 1/ cos +sin^2